A failure of an engine control unit causes it to fail an average of 0.1 times per day. The number of failures is poisson distributed.

TASK

What is the probability that the unit will fail within one week of the previous failure? Round your answer to 4 decimal places. 
Answer #1

Ans:

Given that

Lambda=0.1 per day

=0.7 per week

P(T<=t)=1-exp(-0.7t)

(t is in weeks)

P(T<1 week)=1-exp(-0.7*1)=0.5034